OpenStudy (anonymous):

the abbreviations and units of measurements of: momentum = mass x velocity velocity (speed and direction)

OpenStudy (anonymous):

Momentum=P=kgm/s or Ns Velocity=V=m/s

OpenStudy (anonymous):

Momentum=P=kgm/s or Ns Velocity=V=m/s